OpenStudy (waheguru):

How how to solve this I am confused? please help (-5)^3(-25)^-1 = The answer is 5 but I do not understand why?

OpenStudy (waheguru):

How do we even begin usually we make the same base but we cannot here?

OpenStudy (anonymous):

\[(5)^3(-25)^{-1}\]

OpenStudy (anonymous):

is that you equation

OpenStudy (waheguru):

(-5)^3(-25)^-1 =

OpenStudy (anonymous):

ohh -5 ok

OpenStudy (anonymous):

first then you know what -5^3=?

OpenStudy (waheguru):

I was taught that we can make the same base and then add or subtract the exponents but im confused in this question

OpenStudy (waheguru):

@Reaper534 is that the only way cannoit we make the bases similar in any manner?

OpenStudy (cwrw238):

(-5)^3 = -125 (25)-1 = -1/25

OpenStudy (waheguru):

because when we have variables its much easier to make the same bases/

OpenStudy (cwrw238):